[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Sep 24, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Sep 24 18:38:59 UTC 2016


HSU Shift Log for Sat Sep 24 2016 14:38:53 GMT-0400 (EDT)

Shift: H

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues: None

N-AWIPS/AWIPS2 issues: Reminder to not reboot any NAWIPS workstations this  
weekend.

Other Techincal issues: John noticed that the W: drive is now about 90%  
full (349 GB free out of 3.63 TB).
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160924/36f4874f/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list